### v1

@@ Abstract
The title of a research paper communicates in a succinct style the main theme and, sometimes, the findings of the paper. Coming up with the right title is often an arduous task, and therefore, it would be beneficial to authors if title generation can be automated. In this paper, we fine-tune pre-trained and large language models to generate titles of papers from their abstracts. We also use ChatGPT in a zero-shot setting to generate paper titles. The performance of the models is measured with ROUGE, METEOR, MoverScore, BERTScore and SciBERTScore metrics. Our observations indicate that AI-generated paper titles are generally accurate and appropriate.
@@ 1 Introduction
One of the most important aspects in writing a research paper is coming up with a concise and informative title for the paper. In academic publishing, a title should be not only concise and informative but also engaging, ensuring that authors quickly grasp the paper’s insights. At a first glance, readers should gain a clear understanding of the contributions of the paper. Effective title contain keywords describing the reported research, and sentences in the paper containing words in the title typically reflect the paper’s main theme. Many studies have been devoted to exploring the connection between the title of a paper and its readership as captured by citations and downloads [1, 2, 3]. Short titles reportedly attract more citations [2].
Given the tremendous success of neural language models, especially large language models (LLMs), in various natural language processing (NLP) tasks, it is natural to ask if titles of scientific papers can also be generated by these models. Given the expectation that a title should capture the key import of a paper and yet be a short sequence of words that appeal to a large readership, the task is a challenging one. Title generation can be considered as a special case of abstractive text summarization that aims to distil out the most crucial information form a document.
In this paper, we aim to generate titles of research papers from their abstracts using deep neural models. In particular, we choose several pre-trained transformer models which we fine-tune with a dataset of abstracts and titles. We experiment with both the pre-trained and fine-tuned versions of LLaMA-3 LLM. We also use ChatGPT-3.5 in a zero-shot setting to generate titles from abstracts using prompts. We believe that machine generated titles could be very useful for non-native English speakers who find it difficult to quickly construct a suitable title for their papers, and also for novice researchers making their foray into the field. The generated titles can then be refined by them to more suit their own style and requirement.
The main contributions of this paper are:
@@
    We fine-tuned pre-trained models, namely, T5-base [4], BART-base [5], and PEGASUS-large [6] with CSPubSum dataset to automatically generate titles of research papers. We have also used the open large language model LLaMA-3-8B consist of 8 billion parameters [7][8]. We have contrasted its generation quality with and without fine-tuning it for this task. We have also also used ChatGPT 3.5 in a zero-shot setting to generate title from abstracts.
@@
    We have also curated a dataset of abstracts and titles, called LREC-COLING-2024, on which we evaluate our fine-tuned models without further training.
@@
    We evaluate the performance of the models using ROUGE [9], METEOR [10], MoverScore [11], BERTScore [12] and SciBERTScore metrics. We introduce SciBERTScore in this paper as a variant of BERTScore. We also evaluate the factual accuracy of the model-generated titles at the entity level by employing metrics such as precision-source and F1-target, which were introduced in a study by [13]. Manual evaluation of a small subset of the generated titles has also been carried out.
@@
    Our experiments show that PEGASUS-large outperforms the other models in terms of the chosen metrics. Notably, PEGASUS-large has fewer parameters compared to LLMs like GPT-3.5 and LLaMA.
@@
    A demo that allows the user to choose a pre-trained language model to generate a title from the abstract of a research paper is hosted at https://title-generation-researchpapers.onrender.com/. Additionally, on Hugging Face, we have publicly released the fine-tuned models [^1] https://huggingface.co/TRnlp as well as the LREC-COLING-2024 corpus [^2] https://huggingface.co/datasets/TRnlp/LREC-COLING-2024-Abstract-Title.
@@ 2 Literature survey
The subsequent achievements of sequence-to-sequence models [16], attention-based encoders with beam search and RNN [17], and pointer-generator networks with coverage mechanisms [18] have significant attention to the abstractive text summarization. All are mostly applied on news dataset. The next phase of research in summarization and natural language processing (NLP) in general was propelled by the transformer architecture [19]. It led to the development of various pre-trained language models such as T5 [20], BART [21], and PEGASUS [22] that are initially trained in self-supervised manner on broad, general-purpose text datasets. Their extensive pre-training helps them capture linguistic patterns and knowledge, which can then be fine-tuned for specific tasks across various domains.
In the special case of title generation for a body of text, researchers have proposed methods to generate titles of news articles from their content. Tan et al. [23] proposed a coarse-to-fine approach, which first identifies the important sentences with a with hierarchical attention with a encoder-decoder framework. They tested their model using New York Times (NYT) and DUC-2004 dataset. Our perspective in this paper acknowledges that a title essentially acts as a summary of the content within the research paper [24]. Mishra et al. [25] have proposed a method to generate titles for academic papers. First, they have generated a pool of candidate titles using a pre-trained model, then selected the most suitable one from them, and finally refined it to ensure semantic and syntactic accuracy, thereby enhancing the representativeness of titles for given texts. They have used arXiv [^3] https://tinyurl.com/y9pu6xyp, ACL [26], and ICMLA [27] dataset. Recently, Liu et al. [28] pre-trained a transformer encoder, OAG-BERT, on computer science research papers (including its metadata) and used it to produce paper titles, which were found to be quite acceptable to researchers although extensive evaluation of the generated titles using automatic evaluation measures has not been done. In contrast, we use a diverse set of encoder-decoder and decoder-only transformers in this paper to generate titles from abstracts. We also perform a detailed evaluation of the generated titles. Our work also aligns with the contemporary research [29] [30] that explores the potential of ChatGPT and other artificial intelligence tools in education, research, and scholarly publishing.
@@ 3 Datasets
We have used the CSPubSum dataset provided by Collins et al. [31], which contains URLs of 10147 computer science publications from ScienceDirect [^4] https://www.sciencedirect.com/. We crawled the dataset and organized each example as a pair of abstract and title. The dataset is split into 8120 examples for training, 1014 examples for validation, and 1013 examples for testing. We crawled an additional set of 1000 examples from accepted papers at LREC-COLING 2024555https://aclanthology.org/events/coling-2024/. This dataset, which we hereafter refer to as LREC-COLING-2024, comprises pairs of abstract, title for each paper.
We only train the selected deep neural language models on the train subset of CSPubSum but test them on the test subset of CSPubSum and the corpus of LREC-COLING-2024. While CSPubSum focuses on diverse topics of computer science, they primarily belong to the pre-transformer [19] era since the impact of transformers was produced after 2017 when CSPubSum was released. We wanted to verify how well the trained models perform on a different and more recent dataset on which it is not explicitly trained. So we curated LREC-COLING-2024. The average number of tokens in a title in CSPubSum is 12 while it is 11 for LREC-COLING-2024, although a few titles are longer and some as small as two tokens. A few titles have multiple sentences; a title with two sentences is “Comparative statics effects independent of the utility function. When do we act the same way under risk?”. However, around 82% papers in CSPubSum and 90% papers in LREC-COLING-2024 have no more than 15 tokens in their titles. So when fine-tuning or performing inference with the pre-trained models for title generation, we set the maximum title length to 20 tokens (recollect that due to sub-word generation, token count in a model generally exceeds the raw token count). Another reason for the imposed token limit is that shorter titles are generally more attractive to readers [2].
@@ 4 Methodology
In this section, we detail the fine-tuning process of various pre-trained language models (PLMs) and large language models (LLMs). We have chosen the following models:
@@
    T5-base [4]: It is an encoder-decoder model which is a slight variation of the original Transformer model [19]. Formulating every text processing problem, including translation, question answering, and classification, as a “text-to-text” transformation problem, the same model is trained to perform these diverse tasks. To pre-train the model, random text spans are corrupted/dropped and the model is trained to generate them. T5-base contains 220M parameters. We fine-tune the model with the train subset of CSPubSum.
@@
    BART-base [5]: A denoising autoencoder, it combines Bidirectional and Auto-Regressive Transformers exemplified by BERT [21] and GPT [32], respectively. To train BART, the input text is first corrupted with a noising function, and then the model reconstructs the original text. It is particularly suitable for text generation problems. BART-base is configured with 139M parameters. Like T5-base, it is also fine-tuned on CSPubSum train set.
@@
    PEGASUS-large [6]: It is a Transformer-based encoder-decoder model that is trained on large text collections with an objective function specifically focused on summarization. In particular, pre-training the model involves masking important sentences from an input document and generating them as one output sequence. PEGASUS-large contains 568M parameters. Like the previous two models, we fine-tune PEGASUS-large on training examples of CSPubSum.
@@
    ChatGPT-3.5: We use the instruction and chat-tuned version of the GPT-3 family model, identified as gpt-3.5-turbo-1106, commonly known as ChatGPT 3.5666https://chatgpt.com/. It is a decoder-only model that has been pre-trained on massive text corpora and then further trained with instruction fine-tuning and reinforcement learning with human feedback. It is a large language model with 175B parameters. We use the prompt-based in-context learning setup where we simply prompt the model to generate a title given the abstract.
@@
    LLaMA-3-8B: We use instruction-tuned LLaMA-3-8B model with fine-tuning and without fine-tuning. The LLaMA series of models [7] are decoder-only transformer-based large language models trained exclusively on publicly available datasets (in contrast to the ChatGPT and the GPT series). We fine-tune the LLaMA-3 model (with 8B parameters) on the train subset of CSPubSum.
Note that although the above pre-trained models and LLaMA-3 come in multiple sizes, we choose the smallest sizes that are freely available due to constraints on our computational resources. Although the paper that proposed the PEGASUS model [6] discusses a base and a large model, we could only find the checkpoints for PEGASUS-large in Hugging Face [^7] https://huggingface.co/google/pegasus-large and so, we consider this version. Although GPT-4 is more powerful than GPT-3.5, the former is only available to paid subscribers, and therefore, we do not use it for this research. However, we do expect that, like other deep learning applications, using larger models will enhance the quality of the generated titles.
@@ 4.1 Data processing
We removed only extra spaces from the documents in both datasets. We only retain examples where the abstract length is at least twenty tokens and the paper title length is at least three tokens. Since we impose a limit of 20 tokens on generated title length, we stipulate the abstract to be longer than 20 tokens so that the task can be treated as a text summarization problem.
@@ 4.2 Implementation details
We have chosen the following pre-trained models from the Hugging Face repository for fine-tuning on the CSPubSum dataset: T5-base [^8] https://huggingface.co/t5-base, BART-base [^9] https://huggingface.co/facebook/bart-base, PEGASUS-large [^10] https://huggingface.co/google/pegasus-large . Fine-tuning was performed for 5 epochs. We used batch size of 32, learning rate of 4e-5, and the metric_for_best_model as ROUGE1-F1.
We use the following prompt to generate titles using LLMs, namely ChatGPT-3.5 and LLaMA-3-8B,:
@@
Again during evaluation, both ChatGPT-3.5 and LLaMA-3-8B models were given the above prompt without any in-context examples. Unlike ChatGPT-3.5, LLaMA-3-8B can be trained. But fine-tuning LLaMA-3-8B is an extremely computationally expensive task. Therefore, we use the parameter-efficient fine-tuning technique called Low-Rank Adaptation (LoRA). In particular, we have taken the model from Hugging Face, LLaMA-3-8B  [^11] https://huggingface.co/unsloth/llama-3-8b-bnb-4bit, then loaded it in 4-bit precision to save memory, and finally, fine-tuned it for 5 epochs using a learning rate of 4e-5, train batch size of 32 and eval batch size of 1, rank of the adaptation matrices = r=16 , lora_alpha = 16 (Scaling Factor), and using peft_config for loading LoRA. For fine-tuning and evaluation after fine-tuning, we have used the same prompt that we mentioned above. For comparison, we also include results obtained with LLaMA-3-8B that is not fine-tuned on this dataset, i.e., it is used in zero-shot in-context learning without fine-tuning. We denote this variant as LLaMA-3-8B. In case of ChatGPT-3.5, we use a temperature setting of 0.3, to generate titles based on the abstracts.
For all models, we have set the maximum number of input tokens (i.e., the abstract length) to 512 and output tokens (i.e., the title length) to 20. In all cases, we have measured the memory and compute power consumption using the WandB tool [^12] https://wandb.ai/site. The fine-tuning of the LLaMA-3-8B model took 1 hour 43 minutes for 5 epochs. In the context of pre-trained language models, PEGASUS-large model took 23 minutes, BART-base model took 7 minutes, and T5-base took 13 minutes for 5 epochs of fine-tuning. Training and fine-tuning of the models were done on Tesla A100-SXM4-40GB Colab Pro+ that supports GPU-based training.

### v2

@@ Abstract
The title of a research paper communicates in a succinct style the main theme and, sometimes, the findings of the paper. Coming up with the right title is often an arduous task, and therefore, it would be beneficial to authors if title generation can be automated. In this paper, we fine-tune pre-trained language models to generate titles of papers from their abstracts. Additionally, we use GPT-3.5-turbo in a zero-shot setting to generate paper titles. The performance of the models is measured with ROUGE, METEOR, MoverScore, BERTScore and SciBERTScore metrics. We find that fine-tuned PEGASUS-large outperforms the other models, including fine-tuned LLaMA-3-8B and GPT-3.5-turbo, across most metrics. We also demonstrate that ChatGPT can generate creative titles for papers. Our observations suggest that AI-generated paper titles are generally accurate and appropriate.
@@ 1 Introduction
One of the most important aspects of writing a research paper is coming up with a concise and informative title. In academic publishing, a title should also be engaging, allowing readers to quickly grasp the paper’s key insights. Even at first glance, readers should gain a clear understanding of the paper’s contributions. An effective title includes keywords that describe the research, and sentences in the paper containing words from the title typically reflect its main theme. Many studies have been devoted to exploring the connection between the title of a paper and its readership as captured by citations and downloads [1, 2, 3]. Short titles reportedly attract more citations [2].
Given the tremendous success of neural language models, especially large language models (LLMs), in various natural language processing (NLP) tasks, it is natural to ask if titles of scientific papers can also be generated by these models. Given the expectation that a title should capture the key import of a paper and yet be a short sequence of words that appeal to a large readership, the task is a challenging one. Title generation can be considered as a special case of abstractive text summarization, aiming to distill the most crucial information from a document.
In this paper, we aim to generate titles of research papers from their abstracts using pre-trained language models (PLMs). We fine-tuned several PLMs on a dataset of abstracts and titles. We use the term ‘LLM’ to denote a PLM with more than a billion parameters. We experimented with both the pre-trained and fine-tuned versions of LLaMA-3 LLM. We have also used GPT-3.5-turbo in a zero-shot setting with prompts to generate titles from abstracts. We believe that machine generated titles could be very useful for non-native English speakers who find it difficult to quickly construct a suitable title for their papers, and also for novice researchers making their foray into the field. The generated titles can then be refined by them to better match their personal style and preference. The main contributions of this paper are:
@@
    We fine-tuned the PLMs T5-base [4], BART-base [5], and PEGASUS-large [6] with CSPubSum dataset to automatically generate titles of research papers. We have also used the open-source LLM LLaMA-3-8B [7, 8] and contrasted its generation quality in two settings: with and without fine-tuning it for this specific task. Additionally, we have employed GPT-3.5-turbo in a zero-shot setting to generate titles from abstracts.
@@
    We have curated a dataset of 1000 abstracts and titles, called LREC-COLING-2024, on which we evaluate our fine-tuned models without further training.
@@
    We have measured the performance of the models using ROUGE [9], METEOR [10], MoverScore [11], BERTScore [12] and SciBERTScore metrics. We introduce SciBERTScore in this paper as a variant of BERTScore. We have also evaluated the factual accuracy of the model-generated titles at the entity level by employing metrics such as precision-source and F1-target, which were introduced in a prior study [13]. Manual evaluation of a small subset of the generated titles has also been carried out.
@@
    Our experiments show that PEGASUS-large outperforms the other models in terms of the chosen metrics. Notably, PEGASUS-large has far fewer parameters compared to LLMs like GPT-3.5-turbo and LLaMA.
@@
    We have also analyzed whether ChatGPT can generate creative titles for research papers. In this analysis, we have compared the generated titles with the author-written titles and had them evaluated by human annotators.
  6. 6.
    We have hosted a demo application on Render [^1] https://title-generation-researchpapers.onrender.com/, that allows a user to generate paper titles using pre-trained language models. We have publicly released the fine-tuned models [^2] https://huggingface.co/TRnlp as well as the LREC-COLING-2024 corpus [^3] https://huggingface.co/datasets/TRnlp/LREC-COLING-2024-Abstract-Title on Hugging Face. Additional results are available on our GitHub repository [^4] https://github.com/tohidarehman/Title-Generation-ResearchPapers.
@@ 2 Literature survey
The development of sequence-to-sequence models [16], attention-based encoders with beam search and RNN [17, 18], and pointer-generator networks with coverage mechanisms [19] have enriched the field of abstractive text summarization. These models have primarily been applied to news datasets. The next phase of research in summarization and natural language processing (NLP) in general was propelled by the transformer architecture [20]. It led to the development of various pre-trained language models such as T5 [21], BART [22], and PEGASUS [23], that are initially trained in self-supervised manner on broad, general-purpose text datasets. Their extensive pre-training helps them capture linguistic patterns and knowledge. These pre-trained models can then be fine-tuned for specific tasks across various domains.
In the special case of title generation for a body of text, researchers have proposed methods to generate titles of news articles from their content. Tan et al. [24] proposed a coarse-to-fine approach, which first identifies the important sentences with hierarchical attention in a encoder-decoder framework. They tested their model using New York Times (NYT) and DUC-2004 dataset. Our perspective in this paper acknowledges that a title essentially acts as a summary of the content within the research paper [25]. Mishra et al. [26] have proposed a method to generate titles for academic papers. First, they have generated a pool of candidate titles using a pre-trained model, then selected the most suitable one from them, and finally refined it to ensure semantic and syntactic accuracy, thereby enhancing the representativeness of the titles for the corresponding texts. They have used arXiv [^5] https://tinyurl.com/y9pu6xyp, ACL [27], and ICMLA [28] dataset. More recently, Liu et al. [29] pre-trained a transformer encoder, OAG-BERT, on computer science research papers (including its metadata) and used it to produce paper titles, which were deemed acceptable by researchers. However, comprehensive evaluation of these generated titles using automatic metrics has not been done. Another emerging area of interest is the generation of research highlights from paper abstracts [30, 31, 32, 33]. Distinct from the above, we use a diverse set of encoder-decoder and decoder-only transformers in this paper to generate titles from abstracts. We also perform a detailed evaluation of the generated titles. Our work aligns with the contemporary research that explores the potential of GPT and other artificial intelligence tools in education, research, and scholarly publishing [34, 35].
@@ 3 Datasets
We have used the CSPubSum dataset provided by Collins et al. [36], which contains URLs of 10147 computer science publications from ScienceDirect [^6] https://www.sciencedirect.com/. We crawled the dataset and organized each example as a pair of an abstract and its corresponding title. The dataset is split into 8120 examples for training, 1014 examples for validation, and 1013 examples for testing. We crawled an additional set of 1000 examples from accepted papers at LREC-COLING 2024, that are available on ACL Anthology [^7] https://aclanthology.org/events/coling-2024/ under CC BY 4.0 licence. This dataset, which we hereafter refer to as LREC-COLING-2024, consists of a pair of abstract and title for each paper.
We train the selected language models exclusively on the train subset of CSPubSum but evaluate them both on the test subset of CSPubSum and the corpus of LREC-COLING-2024. While CSPubSum focuses on diverse topics of computer science, they primarily belong to the pre-transformer [20] era as the impact of transformers began to emerge after 2017, the year CSPubSum was released. We wanted to verify how well the trained models perform on a different and more recent dataset that they were not explicitly trained on. This led us to curate LREC-COLING-2024. The average number of tokens in a title in CSPubSum is 12 while it is 11 for LREC-COLING-2024, although a few titles are longer and some as small as two tokens. A few titles have multiple sentences; for example, one title reads: “Comparative statics effects independent of the utility function. When do we act the same way under risk?”. However, around 82% papers in CSPubSum and 90% papers in LREC-COLING-2024 have no more than 15 tokens in their titles. Therefore, when fine-tuning or performing inference with the pre-trained models for title generation, we set the maximum title length to 20 tokens. This limit is also informed by the fact that shorter titles are generally more appealing to readers [2].
@@ 4 Methodology
In this section, we detail the fine-tuning process of various PLMs. We have chosen the following models:
@@
    T5-base [4]: It is an encoder-decoder model which is a slight variation of the original transformer model [20]. Formulating every text processing problem, including translation, question answering, and classification, as a “text-to-text” transformation problem, the same model is trained to perform these diverse tasks. To pre-train the model, random text spans are corrupted/dropped and the model is trained to generate them. T5-base contains 220M parameters.
@@
    BART-base [5]: A denoising autoencoder, it combines bidirectional and auto-regressive transformers exemplified by BERT [22] and GPT [37], respectively. To train BART, the input text is first corrupted with a noising function, and then the model reconstructs the original text. It is particularly suitable for text generation problems. BART-base is configured with 139M parameters.
@@
    PEGASUS-large [6]: It is a transformer-based encoder-decoder model that is trained on large text collections with an objective function specifically focused on summarization. In particular, pre-training the model involves masking important sentences from an input document and generating them as one output sequence. PEGASUS-large contains 568M parameters.
@@
    LLaMA-3-8B: We employed the pre-trained LLaMA-3-8B model [^8] https://ai.meta.com/blog/meta-llama-3/, which consists of 8 billion parameters. We used it in two ways: with fine-tuning and without fine-tuning. The LLaMA series of models [7] are decoder-only Transformer-based large language models trained exclusively on publicly available datasets (in contrast to the GPT series).
@@
    GPT-3.5-turbo: We have used gpt-3.5-turbo-1106999https://platform.openai.com/docs/models/gpt-3-5-turbo. Its web version is commonly known as ChatGPT-3.5101010https://chatgpt.com/. It is a decoder-only model pre-trained on massive text corpora and further fine-tuned using instruction tuning and reinforcement learning from human feedback. It is a successor to GPT-3, a large language model with 175B parameters. We used a prompt-based zero-shot in-context learning setup, where we simply prompted the model to generate a title given the abstract.
We have fine-tuned the first four models on the train subset of CSPubSum. Note that although the above models come in multiple sizes, we have chosen the smallest versions that were freely available at the time of this research, due to constraints on our computational resources. The paper that introduced the PEGASUS model [6] discusses both a base and a large model, but we could only find the checkpoints for PEGASUS-large on Hugging Face and thus, we selected this version. Although GPT-4 is more powerful than GPT-3.5-turbo, the former is only available to paid subscribers, and therefore, we did not use it for this research. However, we anticipate that, as with other deep learning applications, using larger models will improve the quality of the generated titles.
@@ 4.1 Data processing
We removed extra spaces from the documents in both datasets and retained only the examples where the abstract length is at least 20 tokens and the paper title length is at least 3 tokens. Since we imposed a limit of 20 tokens on the generated title length, we stipulated the abstract to be longer than 20 tokens so that the task can be treated as a text summarization problem.
@@ 4.2 Implementation details
We have chosen the following pre-trained models from the Hugging Face repository for fine-tuning on the CSPubSum dataset: T5-base [^11] https://huggingface.co/t5-base, BART-base [^12] https://huggingface.co/facebook/bart-base, PEGASUS-large [^13] https://huggingface.co/google/pegasus-large. Fine-tuning was performed for 5 epochs. We used batch size of 32, learning rate of 4e-5, and the metric_for_best_model as ROUGE1-F1. We used the following prompt to generate titles with GPT-3.5-turbo and LLaMA-3-8B LLMs:
@@
During evaluation, both GPT-3.5-turbo and LLaMA-3-8B models were given the above prompt without any in-context examples. Unlike GPT-3.5-turbo, LLaMA-3-8B can be trained. But fine-tuning LLaMA-3-8B is an extremely computationally intensive task. Therefore, we used the parameter-efficient fine-tuning technique called Low-Rank Adaptation (LoRA). In particular, we have taken the LLaMA-3-8B  [^14] https://huggingface.co/unsloth/llama-3-8b-bnb-4bit model from Hugging Face, then loaded it in 4-bit precision to save memory, and finally, fine-tuned it for 5 epochs using a learning rate of 4e-5, train batch size of 32 and eval batch size of 1, rank of the adaptation matrices r=16 , lora_alpha = 16 (Scaling Factor), and using peft_config for loading LoRA. For both fine-tuning and subsequent evaluation, we utilized the same prompt mentioned earlier. For comparison, we also present results obtained with LLaMA-3-8B that is not fine-tuned on this dataset; we denote this variant as LLaMA-3-8B. In case of GPT-3.5-turbo, we used a temperature setting of 0.3, to generate titles based on the abstracts.
For all models, we have set the maximum number of input tokens (i.e., the abstract length) to 512 and output tokens (i.e., the title length) to 20. In all cases, we have measured the memory and compute power consumption using the WandB tool [^15] https://wandb.ai/site. The fine-tuning of the LLaMA-3-8B model took 1 hour 43 minutes for 5 epochs. In the context of smaller PLMs, for 5 epochs of fine-tuning, PEGASUS-large model took 23 minutes, BART-base model took 7 minutes, and T5-base took 13 minutes. Training and fine-tuning of the models were done on Tesla A100-SXM4-40GB Colab Pro+ that supports GPU-based training.